OpenStudy (sleepyjess):

Decide whether the function is an exponential growth or exponential decay function and find the constant percentage rate of growth or decay. \(f(x) = 7.2 ⋅ 1.08^x\) Exponential growth function; 8% Exponential decay function; 108% Exponential growth function; 108% Exponential growth function; 0.08%

OpenStudy (sleepyjess):

I think it is A but I'm not 100% sure

OpenStudy (sleepyjess):

@amistre64

OpenStudy (amistre64):

A seems fair to me .... but lets dbl chk

OpenStudy (amistre64):

its like a compound interest formula to me

OpenStudy (amistre64):

let x=0 (0,7.2) let x=1 (1,7.776) (7.776-7.2) ---------- = .08, or 8% increase 7.2

OpenStudy (sleepyjess):

ok thank you so much! I was trying to decide between A and D but was leaning mostly towards A